# A microstructural model of transversely isotropic, fibre-reinforced hydrogels

### v1

@@ 5 Modelling unconfined compression experiments
We consider a cylindrical sample of hydrated material of initial radius $R$, height $H$, and porosity $\phi_{0}$. The sample is subjected to an unconfined compression along its longitudinal axis by placing it between two impermeable plates, which are then compressed via an imposed displacement or force $F$, as shown in Fig. 3. During compression, the sample is free to expand in the radial direction. The plates are assumed to be frictionless so that the cylindrical shape of the sample is preserved during compression. Thus, the height of the sample decreases to $h(t)$ and its radius increases to $r_{c}(t)$. The deformation gradient then takes the form
@@ 5.1 Reduction of the governing equations
where $F$ is the applied force. We use the convention that $F<0$ corresponds to axial compression. In this case, (46)–(47) must be simultaneously solved. Once the displacement is found, the radial and circumferential stretches can be computed.
@@ 5.2 Equilibrium response
The equilibrium response is found by setting the time derivatives to zero and seeking a homogeneous deformation with $\beta_{\theta}=\beta_{r}$. The balance of fluid pressure results in $p=0$ everywhere. The referential water fraction can be written in terms of the stretches $\beta_{r}$ and $\beta_{z}$ as $\phi_{w}=\beta_{r}^{2}\beta_{z}-1+\phi_{0}$. For displacement-controlled loading where $\beta_{z}$ is known, the radial stretch $\beta_{r}$ is found by solving $S^{e}_{r}-\beta_{r}\beta_{z}\Pi(\phi_{w})=0$. For force-controlled loading, the radial and axial stretches must be found by simultaneously solving $S^{e}_{r}-\beta_{r}\beta_{z}\Pi(\phi_{w})=0$ and $F=\pi R^{2}[S^{e}_{z}-\beta_{r}^{2}\Pi(\phi_{w})]$.
@@ 5.3 Instantaneous response
The instantaneous response of the sample can be derived by assuming that no fluid is expelled ($\phi_{w}=\phi_{0}$) and that the sample remains homogeneous with $\beta_{\theta}=\beta_{r}=\beta_{z}^{-1/2}$. The fluid in the gel becomes pressurised with a uniform pressure $P>0$ that exceeds the pressure of the surrounding fluid. The balance of total stress at the side of the sample leads to $P=\beta_{z}^{-1/2}S_{r}^{e}-\Pi(\phi_{0})$. If $\beta_{z}$ and hence $\beta_{r}$ and $\beta_{\theta}$ are known due to displacement-controlled loading, then the fluid pressure $P$ can be calculated. However, if $\beta_{z}$ is unknown due to the loading being controlled by force, then the force balance $F=\pi R^{2}(S_{z}^{e}-\beta_{z}^{-3/2}S_{r}^{e})$ must first be solved.
@@ 5.4 Numerical solution
The open-source Python package ucompress.py has been developed to simulate the unconfined compression of nonlinear poroelastic materials using the models developed here. The governing equations are discretised in space using a Chebyshev spectral method and in time using a fully implicit Euler scheme with an analytical Jacobian matrix. The package, along with tutorials, is available at https://github.com/hennessymatt/ucompress.py.
@@ 5.5 Comparison with experiments
Figure 5: Comparison of the model (lines) with experimental (symbols) stress-strain data. Different panels correspond to materials with different nominal fibre fractions $\Phi_{f}$. Solid lines correspond to models that account for fibre recruitment; dashed lines correspond to models without recruitment. Note that positive axial strains in these figures correspond to compressions rather than stretches.
@@
The stress-strain data can be modelled by considering a range of compressive axial stretches $\beta_{z}\leq 1$, corresponding to compressive axial strains of $\epsilon_{z}=1-\beta_{z}$, and computing the axial stress, $S_{z}$, following the procedure outlined in Sec. 5.3. If all fibres in the network are assumed to be mechanically active (no recruitment; $\lambda_{m}=1$), then there are no free parameters in the model. However, the predicted stresses (the dashed lines in Fig. 5) greatly exceed those measured in experiments. This is a strong indication that, despite the materials being pre-loaded due to hydration, some mechanically inactive fibres remain in the network. By accounting for inactive fibres in the model and fitting the maximum recruitment stretch $\lambda_{m}$ to experimental data, excellent agreement is obtained. When fitting $\lambda_{m}$, only one data set for a given fibre fraction is used, yet the model is able to capture both data sets reasonably well. The value of $\lambda_{m}$ obtained by fitting varies with $\Phi_{f}$. We find that for $\Phi_{f}=0.26$, $0.42$, $0.51$ and $0.77$, the best fitting maximum recruitment stretch is $\lambda_{m}=3.0$, $1.8$, $1.7$, and $3.9$, respectively. The non-monotonic relationship between the maximum recruitment stretch and fibre fraction is evident in Fig. 5; the samples with $\Phi_{f}=0.42$ and $0.51$ exhibit a much more rapid increase in stress with strain, characteristic of a stiffer material caused by more fibre engagement.
@@
For the remainder of this section, we focus on FiHy materials that have a nominal fibre fraction $\Phi_{f}=0.42$, as these were subjected to additional experimental testing. By taking the maximum recruitment stretch to be $\lambda_{m}=1.8$, in line with Fig. 5 (b), the mechanical properties of the material are fully characterised. The only remaining fitting parameters are related to the permeability and only influence the time-dependent solutions of the model. After hydration, the height of the sample were $R=5.165$ mm and $h=1.740$ mm, respectively.
The instantaneous and equilibrium response of a FiHy material to different compressive loads was measured. Specifically, the axial stretch $\beta_{z}$ was measured as a function of the applied compressive force $F$. The data is shown in Fig. 6. The model overestimates the magnitude of the instantaneous deformation, but underestimates the equilibrium deformation.
@@
To explore the discrepancy in the predicted equilibrium response, we recall from Sec. 5.2 that, at equilibrium, the applied axial load is balanced by the axial elastic stress and the osmotic stress. For a fixed applied load, $F$, reducing the osmotic stress would increase the axial elastic stress and the axial strain. Since the osmotic stress decreases with porosity, materials with increased hydration will experience greater axial deformation upon application of an applied load. Indeed, by artificially increasing the initial porosity to $\phi_{0}=0.97$ by setting $\chi=0.41$, keeping all other parameters fixed, a much better agreement between the theoretical and experimental equilibrium response is obtained (the dashed-dotted lines in Fig. 6 (b)). As the degree of hydration is determined by a balance between elasticity and osmosis, softer solids tend to become more hydrated. Thus, the need to increase the equilibrium hydration to achieve a reasonable agreement with experimental data suggests that a long-term softening mechanism, such as solid-phase viscoelasticity, could by influencing the mechanics of the material.
@@
Figure 6: Comparison of model predictions (lines) and experimental measurements (symbols) of the (a) instantaneous and (b) equilibrium response of a fibre-reinforced hydrogel to an applied load. We plot the axial stretch $\beta_{z}$ as a function of the applied force $F$. Two values of the initial porosity $\phi_{0}$ are considered.
@@ 6.1 Numerical simulations of the fluid load fraction
The evolution of the fluid load fraction was numerically calculated for a fixed nominal fibre fraction of $\Phi_{f}=0.4$ and a maximum recruitment stretch of $\lambda_{m}=1.5,2.0,3.0$. When a load is applied, the fluid load fraction instantaneously achieves its peak value. Afterwards, it monotonically decays to zero due to poroelastic relaxation; see Fig. 8 (a). More specifically, the application of an applied force increases the fluid pressure above zero, creating a pressure gradient that pushes fluid out of the composite and into the surrounding environment until equilibrium is again established; see Fig. 8 (b).
@@
Numerical calculations show that, as the maximum recruitment stretch increases from $\lambda_{m}=1.5$, corresponding to fibrous networks with more initial slack, the peak fluid load fraction decreases and the duration over which the peak fluid load fraction persists is shortened (Fig. 8 (a)). Thus, removing slack from the fibre network can have a two-fold positive impact on the performance of the material. However, removing slack does accelerate the overall relaxation of the fluid load fraction, causing the new equilibrium state to be reached sooner. This behaviour is expected because the poroelastic relaxation time is inversely proportional to the radial Young’s modulus, which will be larger for materials with fibre networks that are more fully engaged.
@@ 6.2 Insights from the peak fluid load fraction
For the remainder of this section, we focus on the peak fluid load fraction, which can be obtained by solving for the instantaneous response of a sample when a compressive load is applied. From Sec. 5.3, the peak fluid pressure is given by $p=P=\beta_{z}^{-1/2}S_{r}^{e}-\Pi$ and is therefore controlled by two competing mechanisms. Firstly, the radial elastic stresses act to pressurise the fluid. Secondly, the tensile osmotic stress, which provides an additional mechanism to resist axial compression, acts to depressurise the fluid. Thus, maximising the peak fluid load fraction amounts to decreasing the osmotic stress and increasing the radial elastic stress.
@@
The radial elastic stress also plays a key role in controlling the peak fluid load fraction, and is dependent on both the Young’s modulus of the fibres and the maximum recruitment stretch, $\lambda_{m}$. The impact of the fibre modulus on fluid load fraction has been studied by Tan et al. [13]; thus, we focus on the role of fibre engagement. As the maximum recruitment stretch increases, there are more inactive fibres in the network and the radial stiffness decreases. As a result, the fluid becomes less pressurised due to the applied load and the fluid load fraction decreases, as can be seen in Fig. 9. As an extreme case, we have computed the instantaneous response of a fibre-free network. The lack of any radial reinforcement leads to a marked decrease in the fluid load fraction and a loss of fluid pressurisation; see the red diamonds in Fig. 9. Thus, radial reinforcement plays an important role in generating interstitial fluid pressure.
@@ 6.3 Design implications
Moore et al. [16] inferred the fluid load fraction of FiHy samples from unconfined compression experiments. They found that decreasing the nominal fraction of fibres led to a greater fluid load fraction, which disagrees with our findings. However, our analysis of the stress-strain data (see Fig. 5) showed that the amount of slack in the fibre networks varies considerably between samples. Thus, it is likely that the increase in fluid load fraction observed by Moore et al. is due to decreasing the slack in the fibres.
@@ 7 Conclusion
The model implicitly assumes that, whilst their slack is being stretched out, the fibres are able to slip frictionlessly through the hydrogel without mechanical interaction, whereas once fully recruited, they move identically to the matrix. This is a limitation that should be addressed in future work; however, nevertheless, the model is able to capture several key features of fibre-reinforced hydrogels that have been observed experimentally. In particular, the gradual recruitment of fibres allows accurate quantitative agreement with the nonlinear, J-shaped loading curves seen in Fig. 5, as well as qualitative agreement with the time-dependent behaviour displayed in Fig. 7. To improve the quantitative agreement with the transient behaviour, it will likely be necessary to incorporate constitutive viscoelasticity into the fibres themselves. Several non-linear viscoelastic phenomena can be explained by the gradual recruitment of viscoelastic fibres [41].
@@ Appendix B Model derivation using non-equilibrium thermodynamics
By substituting $W=W_{e}(\boldsymbol{\mathsf{F}}_{t})+W_{\text{mix}}(J)$ into (61), using the chain rule and the balance laws (58), and collecting terms leads to
$\left(\frac{\partial W_{e}}{\partial\boldsymbol{\mathsf{F}}_{t}}+J_{t}\frac{\partial W_{\text{mix}}}{\partial J_{t}}\boldsymbol{\mathsf{F}}^{-T}-pJ_{t}\boldsymbol{\mathsf{F}}_{t}^{-T}-\boldsymbol{\mathsf{S}}_{0}\right):\dot{\boldsymbol{\mathsf{F}}}_{t}+\left(p-\mu_{w}\right)\dot{\Phi}_{w}+\nabla_{0}\mu_{w}\cdot\boldsymbol{Q}_{0}\leq 0.$ (62)
@@
$\boldsymbol{\mathsf{S}}_{0}$ $=\frac{\partial W_{e}}{\partial\boldsymbol{\mathsf{F}}_{t}}+J_{t}\frac{\partial W_{\text{mix}}}{\partial J_{t}}\boldsymbol{\mathsf{F}}^{-T}-pJ_{t}\boldsymbol{\mathsf{F}}_{t}^{-T}=\boldsymbol{\mathsf{S}}^{e}_{0}-(\Pi+p)J_{t}\boldsymbol{\mathsf{F}}_{t}^{-T},$ (63b)

### v2

@@ 5 Modelling unconfined compression experiments
We consider a cylindrical sample of hydrated material of initial radius $R$, height $H$, and porosity $\phi_{0}$. The sample is subjected to an unconfined compression along its longitudinal axis by placing it between two impermeable plates, which are then compressed via an imposed displacement or force $F$, as shown in Fig. 3. In this experiment, the sample is surrounded by water to prevent dry out. During compression, the sample is free to expand in the radial direction. The plates are assumed to be frictionless so that the cylindrical shape of the sample is preserved during compression. Thus, the height of the sample decreases to $h(t)$ and its radius increases to $r_{c}(t)$. The deformation gradient then takes the form
@@ 5.1 Reduction of the governing equations
where $F$ is the applied force. We use the convention that $F<0$ corresponds to axial compression. In this case, (46)–(47) must be simultaneously solved. Once the displacement is found, the radial and circumferential stretches, along with the porosity, can be computed.
@@ 5.2 Equilibrium response
The equilibrium response of the material to a load is found by setting the time derivatives to zero and seeking a homogeneous deformation with $\beta_{\theta}=\beta_{r}$. The balance of fluid pressure results in $p=0$ everywhere. The referential water fraction can be written in terms of the stretches $\beta_{r}$ and $\beta_{z}$ as $\phi_{w}=\beta_{r}^{2}\beta_{z}-1+\phi_{0}$. For displacement-controlled loading where $\beta_{z}$ is known, the radial stretch $\beta_{r}$ is found by solving $S^{e}_{r}-\beta_{r}\beta_{z}\Pi(\phi_{w})=0$. For force-controlled loading, the radial and axial stretches must be found by simultaneously solving $S^{e}_{r}-\beta_{r}\beta_{z}\Pi(\phi_{w})=0$ and $F=\pi R^{2}[S^{e}_{z}-\beta_{r}^{2}\Pi(\phi_{w})]$.
@@ 5.3 Instantaneous response
The instantaneous response can be derived by assuming that no fluid is expelled ($\phi_{w}=\phi_{0}$) on short time scales and that the sample remains homogeneous with $\beta_{\theta}=\beta_{r}=\beta_{z}^{-1/2}$. The fluid in the gel becomes pressurised with a uniform pressure $P>0$ that exceeds the pressure of the surrounding fluid. The balance of total stress at the side of the sample leads to $P=\beta_{z}^{-1/2}S_{r}^{e}-\Pi(\phi_{0})$. If $\beta_{z}$ and hence $\beta_{r}$ and $\beta_{\theta}$ are known due to displacement-controlled loading, then the fluid pressure $P$ can be calculated. However, if $\beta_{z}$ is unknown due to the loading being controlled by force, then the force balance $F=\pi R^{2}(S_{z}^{e}-\beta_{z}^{-3/2}S_{r}^{e})$ must first be solved.
@@ 5.4 Numerical solution
The open-source Python package ucompress.py has been developed to provide high-level user-friendly code for simulating the unconfined compression of nonlinear poroelastic materials using the models developed here. The governing equations are discretised in space using a Chebyshev spectral method and in time using a fully implicit Euler scheme with an analytical Jacobian matrix. The package, along with tutorials, is available at https://github.com/hennessymatt/ucompress.py.
@@ 5.5 Comparison with experiments
Figure 5: Comparison of the model (lines) with experimental (symbols) stress-strain data. Different panels correspond to materials with different nominal fibre fractions $\Phi_{f}$. Solid lines correspond to models that account for fibre recruitment; dashed lines correspond to models without recruitment. The maximum recruitment stretch $\lambda_{m}$ for each panel is different; see text for details. Positive axial strains in these figures correspond to compressions rather than stretches.
@@
The stress-strain data can be modelled by considering a range of compressive axial stretches $\beta_{z}\leq 1$, corresponding to compressive axial strains of $\epsilon_{z}=1-\beta_{z}$, and computing the axial stress, $S_{z}$, following the procedure outlined in Sec. 5.3. If all fibres in the network are assumed to be mechanically active (no recruitment; $\lambda_{m}=1$), then there are no free parameters in the model. However, the predicted stresses (the dashed lines in Fig. 5) greatly exceed those measured in experiments. This is a strong indication that, despite the materials being pre-loaded due to hydration, some mechanically inactive fibres remain in the network. By accounting for inactive fibres in the model and fitting the maximum recruitment stretch $\lambda_{m}$ to experimental data, excellent agreement is obtained. When fitting $\lambda_{m}$, only one data set for a given fibre fraction is used, yet the model is able to capture both data sets reasonably well. The value of $\lambda_{m}$ obtained by fitting varies with $\Phi_{f}$. We find that for $\Phi_{f}=0.26$, $0.42$, $0.51$ and $0.77$, the best-fitting maximum recruitment stretch is $\lambda_{m}=3.0$, $1.8$, $1.7$, and $3.9$, respectively. The non-monotonic relationship between the maximum recruitment stretch and fibre fraction is evident in Fig. 5; the samples with $\Phi_{f}=0.42$ and $0.51$ exhibit a much more rapid increase in stress with strain, characteristic of a stiffer material caused by more fibre engagement. The variability in the values of $\lambda_{m}$ obtained from fitting is likely due to material inhomogeneity caused by the electrohydrodynamic fabrication process.
@@
For the remainder of this section, we focus on FiHy materials that have a nominal fibre fraction $\Phi_{f}=0.42$, as these were subjected to additional experimental testing. By taking the maximum recruitment stretch to be $\lambda_{m}=1.8$, in line with Fig. 5 (b), the mechanical properties of the material are fully characterised. The only remaining fitting parameters are related to the permeability and only influence the time-dependent solutions of the model. After hydration, the radius and height of the sample were $R=5.165$ mm and $h=1.740$ mm, respectively.
@@
The instantaneous and equilibrium response of a FiHy material to different compressive loads was measured. Specifically, the axial stretch $\beta_{z}$ was measured as a function of the applied compressive force $F$. The data is shown in Fig. 6. The model overestimates the magnitude of the instantaneous deformation, but underestimates the equilibrium deformation. The latter suggests there is a long-term softening mechanism at play. One explanation for this could be due to an underestimation of the initial degree of hydration. From the equilibrium analysis in Sec. 5.2, we see that the applied axial load is balanced by the axial elastic stress and the osmotic stress. Since the osmotic stress decreases with fluid volume fraction, materials with increased hydration will experience greater axial deformation upon application of a force. Indeed, by setting $\chi=0.41$ to increase the fluid fraction of the initial hydrated state to $\phi_{0}=0.97$, keeping all other parameters fixed, a much better agreement between the theoretical and experimental equilibrium response is obtained (the dashed-dotted lines in Fig. 6 (b)). A second possible mechanism for long-term material softening is solid-phase viscoelasticity, which is expected for fibre-reinforced hydrogels [24] but is not considered in the model here.
@@
Figure 6: Comparison of model predictions (lines) and experimental measurements (symbols) of the (a) instantaneous and (b) equilibrium response of a fibre-reinforced hydrogel to an applied force. We plot the axial stretch $\beta_{z}$ as a function of the applied force $F$. Two values of the initial porosity $\phi_{0}$ are considered.
@@ 6.1 Numerical simulations of the fluid load fraction
The evolution of the fluid load fraction was numerically calculated for a fixed nominal fibre fraction of $\Phi_{f}=0.4$ and a maximum recruitment stretch of $\lambda_{m}=1.5,2.0,3.0$. When a load is applied, the fluid load fraction instantaneously achieves its peak value. Afterwards, it monotonically decays to zero due to poroelastic relaxation; see Fig. 8 (a). More specifically, the application of a force increases the fluid pressure above zero, creating a pressure gradient that pushes fluid out of the composite and into the surrounding environment until equilibrium is again established; see Fig. 8 (b).
@@
Numerical calculations show that, as the maximum recruitment stretch increases from $\lambda_{m}=1.5$, corresponding to fibrous networks with more initial slack, the peak fluid load fraction decreases and the duration over which the peak fluid load fraction persists is shortened (Fig. 8 (a)). Thus, removing slack from the fibre network can have a two-fold positive impact on the performance of the material. However, removing slack does accelerate the overall relaxation of the fluid load fraction, causing the new equilibrium state to be reached sooner. This behaviour is expected because the poroelastic relaxation time is inversely proportional to the radial Young’s modulus, which will be larger for materials with fibre networks that are more engaged.
@@ 6.2 Insights from the peak fluid load fraction
For the remainder of this section, we focus on the peak fluid load fraction, which can be obtained by solving for the instantaneous response of a sample when a compressive load is applied. From Sec. 5.3, the instantaneous fluid pressure is given by $p=P=\beta_{z}^{-1/2}S_{r}^{e}-\Pi$ and is therefore controlled by two competing mechanisms. Firstly, the radial elastic stresses act to pressurise the fluid. Secondly, the tensile osmotic stress, which resists axial compression, acts to depressurise the fluid. Thus, maximising the peak fluid load fraction amounts to decreasing the osmotic stress and increasing the radial elastic stress.
@@
The radial elastic stress also plays a key role in controlling the peak fluid load fraction, and is dependent on both the Young’s modulus of the fibres and the maximum recruitment stretch, $\lambda_{m}$. The impact of the fibre modulus on fluid load fraction has been studied by Tan et al. [13]; thus, we focus on the role of fibre engagement. As the maximum recruitment stretch increases, there are more inactive fibres in the network and the radial stiffness decreases. As a result, the fluid becomes less pressurised due to the applied force and the fluid load fraction decreases, as can be seen in Fig. 9. As an extreme case, we have computed the instantaneous response of a fibre-free network. The lack of any radial reinforcement leads to a marked decrease in the fluid load fraction; see the red diamonds in Fig. 9. Thus, radial reinforcement plays an important role in generating interstitial fluid pressure.
Moore et al. [16] inferred the peak fluid load fraction of FiHy samples from unconfined compression experiments. They found that the fluid load fraction varied non-monotonically with the nominal fibre fraction. Generally, the fluid load fraction was maximised when the (target) ratio of gelatin to PCL was 2:1 by mass. Smaller or larger proportions of gelatin led to a decrease in the fluid load fraction. The results in Fig. 9 indicate that the fluid load fraction is sensitive to the amount of slack in the fibre network, with the fluid load fraction increasing with decreasing slack (or values of $\lambda_{m}$). When fitting the stress-strain data in Fig. 5, we found that materials with intermediate fibre fractions had the least amount of slack and, therefore, would be expected to have the greatest fluid load fractions, which matches with the findings of Moore et al. Thus, the maximum in the fluid load fraction observed by Moore et al. is likely due to variation in the amount of slack in the fibre network across the samples.
@@ 7 Conclusion
The model implicitly assumes that, whilst their slack is being stretched out, the fibres are able to slip frictionlessly through the hydrogel without mechanical interaction, whereas once fully recruited, they move identically to the matrix. This is a limitation that should be addressed in future work. Nevertheless, the model is able to capture several key features of fibre-reinforced hydrogels that have been observed experimentally. In particular, the gradual recruitment of fibres allows accurate quantitative agreement with the nonlinear, J-shaped loading curves seen in Fig. 5, as well as qualitative agreement with the time-dependent behaviour displayed in Fig. 7. To improve the quantitative agreement with the transient behaviour, it will likely be necessary to incorporate constitutive viscoelasticity into the fibres themselves. Several non-linear viscoelastic phenomena can be explained by the gradual recruitment of viscoelastic fibres [41].
@@ Appendix B Model derivation using non-equilibrium thermodynamics
By substituting $W=W_{e}(\boldsymbol{\mathsf{F}}_{t})+W_{\text{mix}}(J_{t})$ into (61), using the chain rule and the balance laws (58), and collecting terms leads to
$\left(\frac{\partial W_{e}}{\partial\boldsymbol{\mathsf{F}}_{t}}+J_{t}\frac{\partial W_{\text{mix}}}{\partial J_{t}}\boldsymbol{\mathsf{F}}_{t}^{-T}-pJ_{t}\boldsymbol{\mathsf{F}}_{t}^{-T}-\boldsymbol{\mathsf{S}}_{0}\right):\dot{\boldsymbol{\mathsf{F}}}_{t}+\left(p-\mu_{w}\right)\dot{\Phi}_{w}+\nabla_{0}\mu_{w}\cdot\boldsymbol{Q}_{0}\leq 0.$ (62)
@@
$\boldsymbol{\mathsf{S}}_{0}$ $=\frac{\partial W_{e}}{\partial\boldsymbol{\mathsf{F}}_{t}}+J_{t}\frac{\partial W_{\text{mix}}}{\partial J_{t}}\boldsymbol{\mathsf{F}}_{t}^{-T}-pJ_{t}\boldsymbol{\mathsf{F}}_{t}^{-T}=\boldsymbol{\mathsf{S}}^{e}_{0}-(\Pi+p)J_{t}\boldsymbol{\mathsf{F}}_{t}^{-T},$ (63b)